OpenStudy (anonymous):

(Solving trig equations) "Find all solutions of: sin^2(x) + 2sin(x)cos(x) - 3cos^2(x) = 0. Hint: Express in terms of tan(x)."

OpenStudy (anonymous):

I don't feel like I'm on the right track with this:

OpenStudy (anonymous):

@pgpilot326 :)

OpenStudy (anonymous):

factor... \[\sin^2x+2\sin x\cos x-3\cos^2 x=0\Rightarrow (\sin x-\cos x)(\sin x + 3\cos x)=0\] set each factor = to 0 and solve.

OpenStudy (anonymous):

that's when you'll use the tangent hint.

OpenStudy (anonymous):

Thanks for your reply, I'll give it a go.

OpenStudy (anonymous):

have fun... gotta run now. don't forget that tangent has a period of pi so add integer multiples of pi to your solutions.

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

I'm struggling to figure out how to write it in terms of tan(x).

OpenStudy (anonymous):

Is this the right thing to do?

OpenStudy (anonymous):

\[\sin x - \cos x = 0 \Rightarrow \sin x = \cos x \Rightarrow \frac{\sin x}{\cos x} = 1 \Rightarrow \tan x = 1\]

OpenStudy (anonymous):

Sweet! I was on the right track with that. I think I've got it now, I'll keep going.

OpenStudy (anonymous):

so what values of x make tan x = 1?

OpenStudy (anonymous):

remember the bit about integer multiples of pi...

OpenStudy (anonymous):

I'm not sure if this is right.

OpenStudy (anonymous):

I've noticed you've been in this thread for a while @ranga , I'm not sure if you're AFK or not, but did I do it right?

OpenStudy (ranga):

sorry. I was answering a couple of other questions and haven't even read this page yet. Are you still looking for an answer?

OpenStudy (anonymous):

Oh, no problem! Yeah, in my second last post I attached an image with my answer but I'm not sure if it's right.

OpenStudy (ranga):

2sin(x) - 3cos(x) = -sin(x)tan(x) divide by -sin(x) -2 + 3cos(x)/sin(x) = tan(x) -2 + 3/tan(x) = tan(x) multiply by tan(x) -2tan(x) + 3 = tan^2(x) tan^2(x) + 2tan(x) - 3 = 0 This is a quadratic equation in tan(x). Solve for tan(x) and then x.

OpenStudy (ranga):

I was starting from the 4th line in your answer.

OpenStudy (anonymous):

Ohh, thank you! I'll give it another go.

OpenStudy (ranga):

alright.

OpenStudy (ranga):

BTW, we can get to the quadratic equation in tan(x) from the given problem quickly as follows: Given problem: sin^2(x) + 2sin(x)cos(x) - 3cos^2(x) = 0 divide throughout by cos^2(x) (assuming cos^2(x) is not zero. When you finally solve for x you should exclude those x that makes cos^2(x) zero). tan^2(x) + 2tan(x) - 3 = 0 solve the quadratic in tan(x) and then x.

OpenStudy (ranga):

tan^2(x) + 2tan(x) - 3 = 0 tan^2(x) + 3tan(x) - tan(x) - 3 = 0 tan(x)(tan(x) + 3) - 1(tan(x) + 3) = 0 (tan(x) - 1)(tan(x) + 3) = 0 tan(x) = 1 or tan(x) = -3 x = arctan(1) or arctan(-3)

OpenStudy (anonymous):

Ohh, interesting. So I did it two different ways: What's the significance of the negative signs?

OpenStudy (ranga):

I will just do the positive part and since trigonometric functions are periodic you can write the solution as (solutions within 2pi) plus/minus 2n(pi) where n = 0, 1, 2, etc.

OpenStudy (anonymous):

Since tan has a period of pi, shouldn't the solutions be +- (pi)n?

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (ranga):

you are welcome.